OpenStudy (anonymous):

Solve x2 + 6x + 7 = 0. x = −1 and x = −5 3 plus or minus square root of 2 negative 3 plus or minus square root of 2 quantity of negative 3 plus or minus square root of 2 all over 2

OpenStudy (anonymous):

is that supposed to be 2x or x square?

OpenStudy (anonymous):

x^2 is x squared 2x is just simply 2x

OpenStudy (anonymous):

its x squared

OpenStudy (anonymous):

Then it should be -1^2+6(-5)+7= 0 1-30+7=0 -29+7=0 -22=0